The voltage in resistor R5 is:

30 R, 21 RS w 512 10 RA ES12V 2012 5,6 V 0,6 V 0.1 v 10V 5,7 V o
Answer #1

Salm Ni 861 RE mon 12v Rişan R₂ RS si yR3 101 ws c B Ry 201 apply KCL at Node A 1,-12 V2 N-V3 + d 6 5 10 V, -120 + 12 V, -12Vapply KCL at Node c V-V2 V2 V₂ - 4 t + 5 2 V₃-2V2 +5V-5V, +V₂ = 0 8V3 – 2V2 -5V, 20. liiij put value of V, from ea. cis to eaV₂=5.709 valt voltage across Ros Vis 11. Vas = V3 -V2 NR = 5-709-5.601 TURS VRg = 0.108 valt CSScanned with CamScanner